OpenStudy (anonymous):

for which equation is x=-1/3 not a solution? a. 7+9x=4 b. 6x-5=7 c. 18+9x=-9 d. 12x+8=4

jhonyy9 (jhonyy9):

a x= -3/9 = -1/3 b x=12/6=2 c x= -27/9 = -3 d x= -4/12 = - 1/3

jhonyy9 (jhonyy9):

- so for b. and c.

OpenStudy (anonymous):

which would be your choice tho?